a financial planner wants to compare the yield of income and growth oriented mutual funds. Fifty thousand dollars is invested in each of a sample of 35 income oriented and 40 growth oriented funds. The mean increase for a two- year period for the income fund is $1100. For the growth oriented funds the mean increase is $1090. At the 0.01 significance level is there a difference in the mean yield of the two funds? Assume that the standard deviation 1=$45 and standard deviation 2 =$55.
I have a very hard time understanding this.
We do not have a textbook for this.
I am to a. State the null and alternative hypothesis
Ho and H1
b. state the decision rule.
c. compute the value of the test statistic
d. Compute the p-value.
e. state the decision of the null hypothesis

Answer by stanbon(48568)

a financial planner wants to compare the yield of income and growth oriented mutual funds. Fifty thousand dollars is invested in each of a sample of 35 income oriented and 40 growth oriented funds. The mean increase for a two- year period for the income fund is $1100. For the growth oriented funds the mean increase is $1090. At the 0.01 significance level is there a difference in the mean yield of the two funds? Assume that the standard deviation 1=$45 and standard deviation 2 =$55.
-----------------------------------------
I have a very hard time understanding this.
We do not have a textbook for this.
I am to:
a. State the null and alternative hypothesis.
You have two samples with mean information so use a 2-sample Z-Test.
Ho: u1-u2 = 0
H1: u1-u2 is not equal to zero
--------------------------
b. state the decision rule.
If the test statistic is lee than -2.576 or greater than 2.576, reject Ho.
----------------------------------
c. compute the value of the test statistic
z(1100-1090) = (1100-1090)/sqrt[(45^2/35) + (55^2/40) = 0.8655
------------------------------------]
d. Compute the p-value.
p-value = 2*P(z > 0.8655)) = 2* = 38.68%
--------------------------------------------
e. state the decision of the null hypothesis
Conclusion: Since p-value is greater than 1%, Fail to reject Ho.
=====================================================================
Cheers,
Stan H.